OpenStudy (thomasfire55):

Find all real and complex roots of the equation below. x^4 - 81 = 0

OpenStudy (thomasfire55):

a) [-9,9,-9i,9i] b) [-3,3] c) [-9,9] d) [-3,3,-3i,3i]

OpenStudy (thomasfire55):

@Mehek14

zepdrix (zepdrix):

Hey Tommy, remember your difference of squares formula? \(\large\rm a^2-b^2=(a-b)(a+b)\) We can apply this rule to 4th powers as well (since 4=2*2) \(\large\rm a^4-b^4=(a^2-b^2)(a^2+b^2)\)

zepdrix (zepdrix):

So we'll start by rewriting 81 as some 4th power, if we're able to. Can you think of a nice way to break down the 81?

zepdrix (zepdrix):

Example: 27=3*3*3 How bout 81?

OpenStudy (thomasfire55):

9?

zepdrix (zepdrix):

Well we could certainly write 81 as 9 squared, \(\large\rm 81=9^2\) But it would be more helpful if we could write it as a fourth power. So let's rewrite our 9 as 3 squared, \(\large\rm \color{orangered}{9}^2=(\color{orangered}{3^2})^2=3^4\) Ok so there is 81 written as a 4th power, it's 3 times 3 times 3 times 3. So our expression looks like this, \(\large\rm x^4-81\) \(\large\rm x^4-3^4\)

zepdrix (zepdrix):

From here we'll apply our difference of squares rule,\[\large\rm x^4-3^4=(x^2-3^2)(x^2+3^2)\]

zepdrix (zepdrix):

This at least gets us closer to our roots.

zepdrix (zepdrix):

Notice what we end up with is another set of `difference of squares`. x^2 minus 3^2. Use the formula again.

zepdrix (zepdrix):

Confused? :O What do you think?

OpenStudy (thomasfire55):

x-3

zepdrix (zepdrix):

Sure, that'll be one of the factors, we should get an x+3 as well though, ya?

zepdrix (zepdrix):

\(\large\rm \color{orangered}{(x^2-3^2)}(x^2+3^2)\) Applying the rule, \(\large\rm \color{orangered}{(x-3)(x+3)}(x^2+3^2)\)

zepdrix (zepdrix):

We're looking for roots, so we'll set our expression equal to 0, \(\large\rm (x-3)(x+3)(x^2+9)=0\) and then apply your `Zero Factor Property`: \(\large\rm (x-3)=0\qquad\qquad(x+3)=0\qquad\qquad (x^2+9)=0\) and solve for x in each case. The last equation should give you 2 roots, since it's a squared x.

OpenStudy (thomasfire55):

so would it be b?

zepdrix (zepdrix):

Maybe :) Why b? I don't like guessers lol :D

OpenStudy (thomasfire55):

cause wouldnt they be 3,-3

zepdrix (zepdrix):

Ok good, 3 from the first factor, -3 from the second factor, what about that last equation? \(\large\rm x^2+9=0\) does this give us any roots?

OpenStudy (thomasfire55):

yes?

zepdrix (zepdrix):

Maybe, solve for x :) Don't be lazy lol :D

OpenStudy (thomasfire55):

3

zepdrix (zepdrix):

Subtract 9 from each side, \(\large\rm x^2=-9\) Square root each side, \(\large\rm x=\pm\sqrt{-9}\) From there, recall what happens when you take the root of a negative number.

OpenStudy (thomasfire55):

so it is b?